OpenStudy (anonymous):

f(x)=In(cos^5(3^4)) find dy/dx

OpenStudy (schrodingers_cat):

Was the three inside the parenthesis supposed to be x?

OpenStudy (anonymous):

f(x)=In(cos^5(3x^4)) find dy/dx

OpenStudy (anonymous):

\[f(x)= \ln (\cos ^{5}(3x^{4})?\]

OpenStudy (schrodingers_cat):

First bring the 5 down in front of the log as that is a property of logs. Then take the derivative of the outside function and continue inward using the chain rule The first step will look like this let u = (cos^5(3x^4)) d/du(5(ln(u)) = 5/u

OpenStudy (schrodingers_cat):

So it will be (5/((cos^5(3x^4)))(d/dx((cos^5(3x^4)))

OpenStudy (schrodingers_cat):

Just keep taking the derivatives using the chain rule :)

OpenStudy (schrodingers_cat):

If I rearrange it like this it is probably easier to read (d/dx((cos^5(3x^4))) (5/((cos^5(3x^4))))

OpenStudy (schrodingers_cat):

Oops with out the cos to the power of five

OpenStudy (schrodingers_cat):

Like this (d/dx((cos(3x^4))) (5/((cos(3x^4))))

OpenStudy (yttrium):

\[f'(x) = \frac{ 1 }{ \cos^5 (3x^4) } (5\cos^4(3x^4)) (12x^3)\] Finally, you'll get \[f'(x) = \frac{ 60x^3 }{ \cos(3x^4) }\] That happened upon applying rules in division of algebraic expression. :)

OpenStudy (schrodingers_cat):

Hmm i got \[-60x^3\sin(3x^4)/(\cos3x^4)\]

OpenStudy (yttrium):

Yeah. I forgot to write about the derivative of cos. It should be \[\frac{ -60x^3\sin3x^4 }{ \cos3x^4 }\] similarly, it will become\[f'(x) = -60x^3\tan3x^4\]

OpenStudy (schrodingers_cat):

Yep :)

OpenStudy (schrodingers_cat):

Do you understand how to get this using the chain rule?

OpenStudy (anonymous):

not really

OpenStudy (schrodingers_cat):

Essentially what you are doing is taking the derivatives the outer functions and working inward.

OpenStudy (schrodingers_cat):

Multiplying the resulting derivatives.

OpenStudy (anonymous):

okay thank you :)

OpenStudy (schrodingers_cat):

Earlier I took the derivative of the outer function which was ln(u) u being (cos^5(3x^4)). Then I substituted back in for u. From the point you just keep peeling away the layers in a way by evaluating (d/dx((cos^5(3x^4))).

OpenStudy (schrodingers_cat):

and so on and so forth until you take the derivative of the innermost function.

OpenStudy (anonymous):

okay thank you :) ... uhh i hate calculus -_-

OpenStudy (schrodingers_cat):

Calculus is great it lets you do so many things!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! :)

OpenStudy (anonymous):

yah but I'm not good at math

OpenStudy (yttrium):

Practice is the best key. :) Remember this, "Geniuses are made, not born." In short, there are no person born to be genius. It's just that they are so exposed into problem solving. :) Just expose yourself there. You'll see.

OpenStudy (anonymous):

aww thank you Yttrium :)

OpenStudy (yttrium):

That's what I believe. :3